OpenStudy (anonymous):

What is the value of y if x = 9? y = 2 + x + 5 A. 7 B. 11 C. 14 D. 16

OpenStudy (anonymous):

Plug it in

OpenStudy (anonymous):

16

OpenStudy (anonymous):

@ItsKansas Please no direct answer. Work them through it. :)

OpenStudy (anonymous):

What is the value of y if x = 9? y = 2 + x + 5 Wherever you see the x, put a 9 in for it. y = 2 + (9) + 5 Now just add them together. 2 + 9 + 5 11 + 5 16 so y = 16

OpenStudy (anonymous):

I was working to it lol

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Do the same thing

OpenStudy (anonymous):

I was working to it

OpenStudy (anonymous):

so a.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what is a's answer?

OpenStudy (anonymous):

a is 23

OpenStudy (anonymous):

OMG I DID IT ON A CALCULATOR ITS A

OpenStudy (anonymous):

ok !

OpenStudy (anonymous):

If k = 5, what is the value of m? m = 4 × 2 + k A. 8 B. 9 C. 13 D. 18

OpenStudy (anonymous):

Do teh same thing...again...

OpenStudy (anonymous):

Whever you see the K, put in a 5 for it